A NEW METHODOLOGY FOR THE DETECTION OF LOW-ABUNDANCE SPECIES IN THE ISM: DETECTION OF INTERSTELLAR CARBODIIMIDE (HNCNH)

\underline{BRETT A. McGUIRE}, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; RYAN A. LOOMIS, Department of Chemistry, University of Virginia, Charlottesville, VA 22904; CAMERON M. CHARNESS, JOANNA F. CORBY, Department of Astronomy, University of Virginia, Charlottesville, VA 22904; GEOFFREY A. BLAKE, Division of Chemistry and Chemical Engineering and Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125; JAN M. HOLLIS, NASA Goddard Space Flight Center, Greenbelt, MD 20771; FRANK J. LOVAS, National Institute of Standards and Technology, Gaithersburg, MD 20899; PHILIP R. JEWELL, and ANTHONY J. REMIJAN, National Radio Astronomy Observatory, Charlottesville, VA 22903.

We present the first interstellar detection of carbodiimide (HNCNH) in observations towards Sgr B2(N) using data from the publicly available Green Bank Telescope PRebiotic Interstellar MOlecular Survey project. Recent laboratory work predicts an abundance of HNCNH of ∼10% of the abundance of its tautomer, cyanamide (NH_{2}CN), or ∼ 2 × 10^{13} cm^{−2} in Sgr B2(N). Given this abundance at LTE conditions, the strongest rotational transitions of HNCNH have intensities at or below the noise level of current observations of this source. A thermal population of HNCNH is therefore likely undetectable. Instead, HNCNH is identified via maser emission features at centimeter wavelengths. This detection presents a new methodology for the detection of low-abundance species and further demonstrates the power of cm-wave observations to make definitive identifications based on a small number of observed features.